OpenStudy (anonymous):

Which equation could be used to solve this problem? The sum of two consecutive integers is 141. Find the integers. A. n + 1= 141 B. 2n + 2 = 141 C. 2n = 141 D. 2n + 1 = 141

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

i'm thinking D @Michele_Laino

OpenStudy (anonymous):

@Haseeb96

OpenStudy (michele_laino):

if n is an integer, then its consecutive is n+1, so we can write this: \[\Large n + \left( {n + 1} \right) = 141\] please simplify

OpenStudy (anonymous):

then A

OpenStudy (michele_laino):

no, since you have to compute this sum: \[\Large n + \left( {n + 1} \right)\]

OpenStudy (anonymous):

then has to be C

OpenStudy (michele_laino):

sorry, C) can't be the right option. Hint: \[\Large n + \left( {n + 1} \right) = n + n + 1\]

OpenStudy (anonymous):

BBBBBBBBBB

OpenStudy (michele_laino):

no, since we have 1 not 2

OpenStudy (anonymous):

whaaaaat. then its D

OpenStudy (michele_laino):

yes! it is D, we can write this: \[\Large n + \left( {n + 1} \right) = n + n + 1 = 2n + 1 = 141\]

OpenStudy (anonymous):

i was thinking D if you read my 2nd comment but thank you @Michele_Laino . :)

OpenStudy (michele_laino):

sorry!

OpenStudy (anonymous):

np. >,<

OpenStudy (michele_laino):

:)